OpenStudy (fanduekisses):

An equation of the line normal to the graph of y = sqrt(3x^2+2x) at (2,4) is ?

OpenStudy (fanduekisses):

\[y = \sqrt{(3x^2+2x)} at (2,4)\]

OpenStudy (fanduekisses):

what does this mean? an equation of the line normal to the graph????

OpenStudy (solomonzelman):

~ find the derivative.

OpenStudy (solomonzelman):

what will the derivative of the function be?

OpenStudy (campbell_st):

the normal is the line that is perpendicular to the tangent at the given point so find the 1st derivative, then the slope of the tangent...

OpenStudy (solomonzelman):

yes.

OpenStudy (fanduekisses):

\[y'= \frac{ 1 }{ 2 }(3x^2+2x)^\frac{ -1 }{ 2 } * 6x+2\]

OpenStudy (solomonzelman):

you mean (6x+2) like this?

OpenStudy (solomonzelman):

then it is correct.

OpenStudy (solomonzelman):

you can simplify it a bit though.

OpenStudy (solomonzelman):

\(\large\color{black}{ f'(x)=\frac{\LARGE 1 }{\LARGE 2}\cdot (\sqrt{3x^2+2x})^{-1/2}\cdot(6x+2) }\) please re-write this for me, with a positive exponent, and most reduced terms.

OpenStudy (solomonzelman):

lost?

OpenStudy (fanduekisses):

\[\frac{ 1 }{ \sqrt{3x^2+2x} }\times3x+1\]

OpenStudy (fanduekisses):

Is it like that?

OpenStudy (freckles):

\[f'(x)=\frac{3x+1}{\sqrt{3x^2+2x}}\]

OpenStudy (freckles):

looks good

OpenStudy (fanduekisses):

:D

OpenStudy (fanduekisses):

now what happens?

OpenStudy (freckles):

well do you know how to find the slope of the tangent line to f at (2,4)

OpenStudy (fanduekisses):

I know how to find slope between two points, but not of a tangent line (or don;t remember :( )

OpenStudy (freckles):

well f' gives you the slope number for each point (x,y) on f

OpenStudy (freckles):

evaluate f' at (2,4) to get the slope of the tangent line to f at (2,4)

OpenStudy (freckles):

we will then find the slope of line that runs perpendicular to that tangent line by flipping your number and taking the opposite of it

OpenStudy (fanduekisses):

OK lol Yeaahhh the derivative is the rate of change ok so I'll plug in

OpenStudy (freckles):

yea

OpenStudy (fanduekisses):

I get 1.75

OpenStudy (freckles):

your normal line will be in this form \[y-y_1=\frac{-1}{m}(x-x_1)\] where m was the slope of the tangent line

OpenStudy (fanduekisses):

\[y-4=-\frac{ 4 }{ 7 }(x-2)\]

OpenStudy (freckles):

\[f'(2)=\frac{3(2)+1}{\sqrt{3(2)^2+2(2)}}=\frac{7}{4}\]

OpenStudy (freckles):

looks great

OpenStudy (freckles):

you are done unless you want the line in slope-intercept form

OpenStudy (fanduekisses):

y=-4/7x+36/7

OpenStudy (freckles):

\[\text{checking...} \\ y=\frac{-4}{7}(x-2)+4 \\ y=\frac{-4}{7}x+\frac{8}{7}+4 \\ y=\frac{-4}{7}x+\frac{8+28}{7}\] yep looks great also

OpenStudy (freckles):

on your previous question on the calculator I'm kinda thinking you entered in 100/20 sqrt(3) instead of 100/(20 sqrt(3)) into calculator and tht is why you got 5sqrt(3) instead of 5/sqrt(3)

OpenStudy (fanduekisses):

yeaaahh that is what I did lol

OpenStudy (solomonzelman):

VERIFICATION: https://www.desmos.com/calculator/os0v8fyubn

OpenStudy (solomonzelman):

you can use ~ for a space, btw.

OpenStudy (solomonzelman):

more space? ~~ Even more ~~~ and on....

OpenStudy (fanduekisses):

thanks @freckles and @SolomonZelman